OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (14, -2) and is perpendicular to 7x – 10y = 18 y=-7/10x+9/5 is where i am stuck at?

OpenStudy (anonymous):

okay coz its perpendicular then the slope of the other line is 10/7

OpenStudy (anonymous):

so i need to y=mx+b

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

okay two lines if they are perpendicular, there slope if multiply together yu get -1

OpenStudy (anonymous):

so coz the slope of the first equation was -7/10 then yu take the reciprocal of -7/10 and multiply by -1

OpenStudy (anonymous):

y=(10/7)x+b

OpenStudy (anonymous):

and then use the given point (14,-2) -2=20+b b=-22 y=(10/7)x-22

OpenStudy (radar):

That slope might be (-10/7) for the newly found perpendicular line. Double check!

OpenStudy (anonymous):

yes thats true..... i jst kept on working from where she got stuck

OpenStudy (radar):

miss_texas are you still there?

OpenStudy (anonymous):

yes i am getting a diff answer

OpenStudy (anonymous):

y=10/7x+9/5 (y-y1)=m(x-x1) (y+2)=10/7(x-14)

OpenStudy (radar):

First express the given line in the slope intercept form y=mx+b. Like this: 7x-10y=18 -10y=-7x+18 y=(-7/-10)x +18 or y=7x/10+18 Slope is 7/10 For a line to be perpendicular to this line, its slope must be the negative reciprocal as jkip706 explained. Or the slope must be -10/7. now you can plug one of the points in and create the equation for the new line. the point P is (14,-2) x=14, y=-2 -2=(-10/7)(14)+b -2=-20+b b=-2+20=18 thats all you need now write the equation: y=(-10/7)x+18

OpenStudy (anonymous):

is that suppose to be my final answer

OpenStudy (radar):

Yes, and where you got stuck was your treatment to the signs. If you had treated the signs correctly you would not come up with y=(-7/10)x+9/5 but would of gotten y=(7/10)x-9/5

OpenStudy (radar):

Did you find your error?